1.  H&M has been selling 15,000 jackets per month for 355 ETB. When they increased the price to 500 ETB, they sold only 12,000 Jackets. What is the demand elasticity? If the marginal cost is 250 per Jacket, what will be the desired markup price? Was raising the price profitable?


1
Expert's answer
2022-01-31T07:39:04-0500

Solution:

Demand elasticity = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


Q1 = 15,000                                     P1 = 355

Q2 = 12,000                                     P2 = 500

 

Demand elasticity = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"


= "\frac{12,000 -15,000}{(12,000+15,000)/2 } \div \frac{500 -355}{(500+355)/2 } = \frac{-3000}{13,500} \div\frac{145}{427.5} = \frac{-0.22}{0.34} = -0.65"


Demand elasticity = 0.65

 

Desired markup price = "\frac{(selling\; price - marginal \;cost)}{marginal\; cost} \times 100\% = \frac{(500 - 250)}{250} \times 100\% = 100\%"

 

Raising the price was profitable since the demand is price inelastic and the H&M was able to double the profits.